# Test-Time Training on Video Streams

### v1

@@ 5.1 Additional Baselines
Alternative architectures. The authors of Mask2Former did not evaluated it on KITTI-STEP. We benchmark Mask2Former on the KITTI-STEP validation set against two other popular models of comparable size: SegFormer B4 [74] (64.1M), and DeepLabV3+/RN101 [9] (62.7M), which is used by [69]. Mean IoU are, respectively, 42.0% and 53.1%. Given Main Task Only in Table 1 has 53.8%, we can verify that our pre-trained model (69M) is indeed the state-of-the-art on KITTI-STEP. For COCO segmentation, the authors of Mask2Former have already compared with alternative architectures [10], so we do not repeat their experiments.
@@
Alternative techniques for TTT. Self-supervision with MAE is only one particular technique for test-time training. Subsection 4.2 describes an outer loop, and any technique that does not use ground truth labels can be used to update the model inside the loop. We experiment with three most promising ones according to prior work: self-training [69], layer norm (LN) adaptation [54], and Tent [70]. For self-training, our implementation significantly improves on the version in [69]. Please refer to Appendix A for an in depth discussion of these three techniques.
Class balancing. Volpi et al. propose a heuristic that is applicable when implicit memory is used [69]: Record the number of predicted classes, for the initial model $h\circ f_{0}$ and the current model $h\circ f_{t}$. Reset the model parameters when the difference is large enough, in which case the predictions of the current model have likely collapsed. To compare with [69], we evaluate this heuristic on self-training and Tent. Methods with class balancing are appended with Class. It cannot be applied to LN Adapt, which does not actually modify the trainable parameters in the model.
@@
CityScapes-VPS [32] 1.8 3000 17 19
DAVIS [49] 3.5 3455 30 -
YouTube-VOS [76] 4.5 123,467 30 94
KITTI-STEP [72] 40 8,008 10 19
@@ 5.2 Semantic Segmentation on KITTI-STEP
KITTI-STEP [72] contains 9 validation videos and 12 test videos of urban driving scenes.222 KITTI-STEP is originally designed to benchmark instance-level tracking, and has a separate test set held-out by the organizers. The official website evaluates only tracking-related metrics on this test set. Therefore, we perform our own evaluation using the segmentation labels. Since we do not perform regular training on KITTI-STEP, we use the training set as test set. At the rate of 10 frames-per-second, these videos are the longest – up to 106 seconds – among public datasets with dense pixel-wise annotations. All hyper-parameters, even for COCO Videos, are selected on the KITTI-STEP validation set. Joint training is performed on CityScapes [12], another driving dataset with exactly the same 19 categories as KITTI-STEP, but containing still images instead of videos.
@@
Baseline techniques that adapt the normalization layers alone do not help at all in these evaluations. This agrees with the evidence in [69]: LN Adapt and Tent help significantly on datasets with synthetic corruptions, but do not help on the real-world dataset (CityScapes).
TTT-MAE optimizes for only 1 iteration per frame, and is 2x slower than the baseline. Comparing with TTT-MAE [19], which optimizes for 20 iterations per image, our method runs much faster. Again, this is because our implicit memory takes advantage of temporal smoothness to get a better initialization for every frame. Resetting parameters is wasteful on videos, because the adjacent frames are very similar.
@@ 5.3 COCO Videos
We collected 10 videos, each about 5 minutes, annotated by professionals, in the same format as for COCO instance and panoptic segmentation [38]. The benchmark metrics are also the same as in COCO: average precision (AP) for instance and panoptic quality (PQ) for panoptic. To put things into perspective, each of the 10 videos alone contains more frames, at the same rate, than all of the videos combined in the KITTI-STEP validation set. We compare this new dataset with other publicly available ones in Table 2.
@@
All videos are egocentric, similar to the visual experience of a human walking around. In particular, they do not follow any tracked object like in Oxford Long-Term Tracking [64] or ImageNet-Vid [55]. Objects leave and enter the camera’s view all the time. Unlike KITTI-STEP and CityScapes that focus on self-driving scenes, our videos are both indoors and outdoors, taken from diverse locations such as sidewalks, markets, schools, offices, restaurants, parks and households. Figure 4 shows random frames from COCO Videos and their ground truth labels.
@@
Zhang et al. [80] 62.39 5.00 $\pm$ 0.19 0.180 22.27 0.924
Main Task Only [10] 59.96 5.23 $\pm$ 0.12 0.216 20.42 0.881
@@ 5.4 Video Colorization
The goal of colorization is to add realistic RGB colors to gray-scale images  [35, 78]. Our goal here is to demonstrate the generality of online TTT-MAE, not to achieve the state-of-the-art.
Following [80], we simply treat colorization as a supervised learning problem. We use the same architecture as for segmentation – Swin Transformer with two heads, trained on ImageNet [14] to predict the colors given input images processed as gray-scale. We only make the necessary changes to map to a different output space, and do not use domain-specific techniques, e.g., perceptual losses, adversarial learning, and diffusion models. Our bare-minimal baseline already achieves results comparable, if not superior, to those in [80]. Our method uses exactly the same hyper-parameters as for segmentation. All of our colorization experiments were completed in a single run.
@@
Our method outperforms the baseline and [80] on all metrics except SSIM. It is a field consensus that PSNR and SSIM often misrepresent actual visual quality because colorization is inherently multi-modal [80, 81], but we still include them for completeness. Please see the project website for the complete set of the original and colorized videos. Our method visually improves the quality in all of them comparing to the baseline, especially in terms of consistency across frames.
@@ 6.1 Empirical Analysis
Table 4 contains ablations on our two forms of memory: implicit and explicit. Both forms of memory contribute to the improved performance of our final method over single-image TTT-MAE [19]. Beyond these basic ablations, we further analyze three aspects of our method.
@@ 6.2 Theoretical Analysis
Remark on assumptions. Assumption 1, that neural networks are strongly convex around their local minima, is widely accepted in the deep learning theory community [2, 83, 71]. Assumption 2 is simply temporal smoothness in L2 norm; any norm could be used here as long as the norm in Assumption 1 for strong convexity is also changed accordingly. Assumption 3, that the main task and self-supervised task have correlated gradients, comes from the theoretical analysis of [61].
@@ 7 Discussion
Test-time training on nearest neighbors (TTT-NN). Here is an alternative heuristic for test-time training: For each test instance, retrieve its nearest neighbors from the training set, and fine-tune the model on those neighbors before applying it to the test instance. As discussed in Subsection 2.2, this simple heuristic has been shown to significantly improve prediction quality in [7] three decades ago, and more recently in [25] for large language models. Given temporal smoothness, our sliding window can be seen as retrieving neighbors of the current frame. The only difference is that our method retrieves from past unlabeled test instances, while TTT-NN retrieves from labeled training data. Therefore, our method must use self-supervision for TTT, instead of supervised learning on the neighbors as in TTT-NN. But this difference also implies an important advantage for our method: Given temporal smoothness, our neighbors will always be actually relevant, while the neighbors from the training set might not.
@@ A.1 Self-Training
Self-training is a popular technique in semi-supervised learning [51, 52, 85, 4, 3] and domain adaptation [34, 86, 44, 39, 59]. It is also evaluated in [69], but is shown to produce inferior performance. We experiment with both its original form, and our own designs that actually improve this baseline.
@@
Our first design improvement is incorporating the confidence threshold $\lambda$. In [69], all predictions are pseudo-labels, regardless of confidence. Experiments show that for low $\lambda$, or with $\lambda=0$ in [69], self-training is noisy and unstable, as expected.
However, for high $\lambda$, there is limited learning signal, e.g. little gradient, since $f$ is already very confident about the pseudo-label. Our second design improvement, inspired by [58], is to make learning more challenging with an already confident prediction, by masking image patches in $x$. In [58], masking is applied sparingly on 2.5% of the pixels in average. We mask 80% of the pixels, inspired by [27].
@@ A.2 Layer Norm Adapt
Prior work [54] shows that simply recalculating the batch normalization (BN) statistics works well for unsupervised domain adaptation. [69] applies this technique to video streams by accumulating the statistics with a forward pass on each frame once it is revealed. Since modern transformers use layer normalization (LN) instead, we apply the same technique to LN.
@@ A.3 Tent
The normalization layers (BN and LN) also contain trainable parameters that modify the statistics. Optimizing those parameters requires a self-supervised objective. Tent [70] is an objective for learning only those parameters at test time, by minimizing the softmax entropy of the predicted distribution over classes. We update the LN statistics and parameters with Tent, in the same loop as our method, also using implicit and explicit memory. Hyper-parameters are searched on the KITTI-STEP validation set to be optimal for Tent.
Figure 9: Samples results for video colorization on the Lumiere Brothers films. Top: Using Zhang et al. [80]. Middle: Using our own baseline, Mask2Former with Main Task Only, which is already comparable, if not superior to [80]. Bottom: After applying online TTT-MAE on the baseline. Our colors are more vibrant and consistent within regions.
@@ Appendix C Proof of Theorem 1
Proof of lemma. It is a well known fact in convex optimization [8] that for $f$ $\alpha$-strongly convex and continuously differentiable,

### v3

@@ 5.1 Additional Baselines
Alternative architectures. The authors of Mask2Former did not evaluated it on KITTI-STEP. We benchmark Mask2Former on the KITTI-STEP validation set against two other popular models of comparable size: SegFormer B4 [xie2021segformer] (64.1M), and DeepLabV3+/RN101 [chen2017rethinking] (62.7M), which is used by [volpi2022road]. Mean IoU are, respectively, 42.0% and 53.1%. Given Main Task Only in Table 1 has 53.8%, we can verify that our pre-trained model (69M) is indeed the state-of-the-art on KITTI-STEP. For COCO segmentation, the authors of Mask2Former have already compared with alternative architectures [cheng2021masked], so we do not repeat their experiments.
@@
Alternative techniques for TTT. Self-supervision with MAE is only one particular technique for test-time training. Subsection 4.2 describes an outer loop, and any technique that does not use ground truth labels can be used to update the model inside the loop. We experiment with three most promising ones according to prior work: self-training [volpi2022road], layer norm (LN) adaptation [schneider2020improving], and Tent [wang2020tent]. For self-training, our implementation significantly improves on the version in [volpi2022road]. Please refer to Appendix A for an in depth discussion of these three techniques.
Class balancing. Volpi et al. propose a heuristic that is applicable when implicit memory is used [volpi2022road]: Record the number of predicted classes, for the initial model $h\circ f_{0}$ and the current model $h\circ f_{t}$. Reset the model parameters when the difference is large enough, in which case the predictions of the current model have likely collapsed. To compare with [volpi2022road], we evaluate this heuristic on self-training and Tent. Methods with class balancing are appended with Class. It cannot be applied to LN Adapt, which does not actually modify the trainable parameters in the model.
@@
CityScapes-VPS [kim2020video] 1.8 3000 17 19
DAVIS [Pont-Tuset_arXiv_2017] 3.5 3455 30 -
YouTube-VOS [xu2018youtube] 4.5 123,467 30 94
KITTI-STEP [weber2021step] 40 8,008 10 19
@@ 5.2 Semantic Segmentation on KITTI-STEP
KITTI-STEP [weber2021step] contains 9 validation videos and 12 test videos of urban driving scenes.222 KITTI-STEP is originally designed to benchmark instance-level tracking, and has a separate test set held-out by the organizers. The official website evaluates only tracking-related metrics on this test set. Therefore, we perform our own evaluation using the segmentation labels. Since we do not perform regular training on KITTI-STEP, we use the training set as test set. At the rate of 10 frames-per-second, these videos are the longest – up to 106 seconds – among public datasets with dense pixel-wise annotations. All hyper-parameters, even for COCO Videos, are selected on the KITTI-STEP validation set. Joint training is performed on CityScapes [cordts2016cityscapes], another driving dataset with exactly the same 19 categories as KITTI-STEP, but containing still images instead of videos.
@@
Baseline techniques that adapt the normalization layers alone do not help at all in these evaluations. This agrees with the evidence in [volpi2022road]: LN Adapt and Tent help significantly on datasets with synthetic corruptions, but do not help on the real-world dataset (CityScapes).
TTT-MAE optimizes for only 1 iteration per frame, and is 2x slower than the baseline. Comparing with TTT-MAE [ttt-mae], which optimizes for 20 iterations per image, our method runs much faster. Again, this is because our implicit memory takes advantage of temporal smoothness to get a better initialization for every frame. Resetting parameters is wasteful on videos, because the adjacent frames are very similar.
@@ 5.3 COCO Videos
We collected 10 videos, each about 5 minutes, annotated by professionals, in the same format as for COCO instance and panoptic segmentation [lin2014microsoft]. The benchmark metrics are also the same as in COCO: average precision (AP) for instance and panoptic quality (PQ) for panoptic. To put things into perspective, each of the 10 videos alone contains more frames, at the same rate, than all of the videos combined in the KITTI-STEP validation set. We compare this new dataset with other publicly available ones in Table 2.
@@
All videos are egocentric, similar to the visual experience of a human walking around. In particular, they do not follow any tracked object like in Oxford Long-Term Tracking [valmadre2018long] or ImageNet-Vid [shankar2021image]. Objects leave and enter the camera’s view all the time. Unlike KITTI-STEP and CityScapes that focus on self-driving scenes, our videos are both indoors and outdoors, taken from diverse locations such as sidewalks, markets, schools, offices, restaurants, parks and households. Figure 4 shows random frames from COCO Videos and their ground truth labels.
@@
Zhang et al. [zhang2016colorful] 62.39 5.00 $\pm$ 0.19 0.180 22.27 0.924
Main Task Only [cheng2021masked] 59.96 5.23 $\pm$ 0.12 0.216 20.42 0.881
@@ 5.4 Video Colorization
The goal of colorization is to add realistic RGB colors to gray-scale images  [lei2019fully, zhang2019deep]. Our goal here is to demonstrate the generality of online TTT-MAE, not to achieve the state-of-the-art.
Following [zhang2016colorful], we simply treat colorization as a supervised learning problem. We use the same architecture as for segmentation – Swin Transformer with two heads, trained on ImageNet [deng2009imagenet] to predict the colors given input images processed as gray-scale. We only make the necessary changes to map to a different output space, and do not use domain-specific techniques, e.g., perceptual losses, adversarial learning, and diffusion models. Our bare-minimal baseline already achieves results comparable, if not superior, to those in [zhang2016colorful]. Our method uses exactly the same hyper-parameters as for segmentation. All of our colorization experiments were completed in a single run.
@@
Our method outperforms the baseline and [zhang2016colorful] on all metrics except SSIM. It is a field consensus that PSNR and SSIM often misrepresent actual visual quality because colorization is inherently multi-modal [zhang2016colorful, zhang2017real], but we still include them for completeness. Please see the project website for the complete set of the original and colorized videos. Our method visually improves the quality in all of them comparing to the baseline, especially in terms of consistency across frames.
@@ 6.1 Empirical Analysis
Table 4 contains ablations on our two forms of memory: implicit and explicit. Both forms of memory contribute to the improved performance of our final method over single-image TTT-MAE [ttt-mae]. Beyond these basic ablations, we further analyze three aspects of our method.
@@ 6.2 Theoretical Analysis
Remark on assumptions. Assumption 1, that neural networks are strongly convex around their local minima, is widely accepted in the deep learning theory community [pmlr-v97-allen-zhu19a, zhong2017recovery, wang2021hidden]. Assumption 2 is simply temporal smoothness in L2 norm; any norm could be used here as long as the norm in Assumption 1 for strong convexity is also changed accordingly. Assumption 3, that the main task and self-supervised task have correlated gradients, comes from the theoretical analysis of [sun2020test].
@@ 7 Discussion
Test-time training on nearest neighbors (TTT-NN). Here is an alternative heuristic for test-time training: For each test instance, retrieve its nearest neighbors from the training set, and fine-tune the model on those neighbors before applying it to the test instance. As discussed in Subsection 2.2, this simple heuristic has been shown to significantly improve prediction quality in [bottou1992local] three decades ago, and more recently in [hardt2023test] for large language models. Given temporal smoothness, our sliding window can be seen as retrieving neighbors of the current frame. The only difference is that our method retrieves from past unlabeled test instances, while TTT-NN retrieves from labeled training data. Therefore, our method must use self-supervision for TTT, instead of supervised learning on the neighbors as in TTT-NN. But this difference also implies an important advantage for our method: Given temporal smoothness, our neighbors will always be actually relevant, while the neighbors from the training set might not.
@@ Acknowledgements
Figure 7: Semantic segmentation predictions for adjacent frames from a video in KITTI-STEP. Top: Results using a fixed model baseline without TTT. Predictions are inconsistent between the two frames. The terrain on the right side of the road is incompletely segmented in both frames, and the terrain on the left is incorrectly classified as a wall on the first frame. Bottom: Results using online TTT-MAE, by the same model, on the same frames as top. Predictions are now consistent and correct.
Figure 8: Panoptic segmentation predictions for adjacent frames from a video in our new COCO Videos dataset. Top: Results using a fixed model baseline without TTT. Predictions are inconsistent between the two frames. Bottom: Results using online TTT-MAE, by the same model, on the same frames as top. Predictions are now consistent and correct. Please zoom in to see the instance labels.
@@ A.1 Self-Training
Self-training is a popular technique in semi-supervised learning [Radosavovic_2018_CVPR, rosenberg2005semi, zoph2020rethinking, asano2019self, asano2020labelling] and domain adaptation [kumar2020understanding, zou2018unsupervised, mei2020instance, liu2021energy, spadotto2021unsupervised]. It is also evaluated in [volpi2022road], but is shown to produce inferior performance. We experiment with both its original form, and our own designs that actually improve this baseline.
@@
Our first design improvement is incorporating the confidence threshold $\lambda$. In [volpi2022road], all predictions are pseudo-labels, regardless of confidence. Experiments show that for low $\lambda$, or with $\lambda=0$ in [volpi2022road], self-training is noisy and unstable, as expected.
However, for high $\lambda$, there is limited learning signal, e.g. little gradient, since $f$ is already very confident about the pseudo-label. Our second design improvement, inspired by [sohn2020simple], is to make learning more challenging with an already confident prediction, by masking image patches in $x$. In [sohn2020simple], masking is applied sparingly on 2.5% of the pixels in average. We mask 80% of the pixels, inspired by [mae].
@@ A.2 Layer Norm Adapt
Prior work [schneider2020improving] shows that simply recalculating the batch normalization (BN) statistics works well for unsupervised domain adaptation. [volpi2022road] applies this technique to video streams by accumulating the statistics with a forward pass on each frame once it is revealed. Since modern transformers use layer normalization (LN) instead, we apply the same technique to LN.
@@ A.3 Tent
The normalization layers (BN and LN) also contain trainable parameters that modify the statistics. Optimizing those parameters requires a self-supervised objective. Tent [wang2020tent] is an objective for learning only those parameters at test time, by minimizing the softmax entropy of the predicted distribution over classes. We update the LN statistics and parameters with Tent, in the same loop as our method, also using implicit and explicit memory. Hyper-parameters are searched on the KITTI-STEP validation set to be optimal for Tent.
Figure 9: Samples results for video colorization on the Lumiere Brothers films. Top: Using Zhang et al. [zhang2016colorful]. Middle: Using our own baseline, Mask2Former with Main Task Only, which is already comparable, if not superior to [zhang2016colorful]. Bottom: After applying online TTT-MAE on the baseline. Our colors are more vibrant and consistent within regions.
@@ Appendix C Proof of Theorem 1
Proof of lemma. It is a well known fact in convex optimization [bubeck2015convex] that for $f$ $\alpha$-strongly convex and continuously differentiable,